Rohan deposit 25000 into account. He deposited some amount in saving account at simple interest 6% p.a. remaining amount in fixed deposit @12% p.a. if he gets 5700 at total interest after two years then the amount he deposit in fixed deposit is?
Answer by mananth(16949)

Part I 6.00% per annum ------------- Amount invested =x
Part II 12.00% per annum ------------ Amount invested = y =
25000
Interest----- 2850

Part I 6.00% per annum ---x
Part II 12.00% per annum ---y
Total investment
x + 1 y= 25000 -------------1
Interest on both investments
for one year interest is 2850
6.00% x + 12.00% y= 2850
Multiply by 100
6 x + 12 y= 285000.00 --------2
Multiply (1) by -6
we get
-6 x -6 y= -150000.00
Add this to (2)
0 x 6 y= 135000
divide by 6
y = 22500
Part I 6.00% $ 2500
Part II 12.00% $ 22500

CHECK
2500 --------- 6.00% ------- 150.00
22500 ------------- 12.00% ------- 2700.00
Total -------------------- 2850.00

m.ananth@hotmail.ca